### v1

# Mathematical Anatomy of Neutrino Decoherence in Red Turbulence: A Fractional Calculus Approach

Abstract
We develop an exact framework for neutrino decoherence in power-law correlated turbulent matter, as encountered in core-collapse supernovae. Employing the Nakajima–Zwanzig projection technique, we derive an exact non-Markovian master equation for the neutrino density matrix. For red turbulence spectra, the memory kernel follows $K(t)\propto t^{-\nu}$, leading naturally to a fractional calculus description. To maintain mathematical rigor for $\nu\geq 1$, we introduce an exponential cutoff at the dissipation scale, regularizing ultraviolet divergences.
The exact Laplace-space solution for the survival probability is obtained. In the high-density matter basis relevant to supernovae, the solution is expressed through Mittag-Leffler functions, establishing a direct link to anomalous diffusion phenomena. For extremely red spectra ($\nu<0$), the memory integral corresponds to a higher-order fractional operator.
Our work clarifies how spectral index, regularization scale, and decoherence efficiency interrelate, providing a complete analytical description and practical tools for supernova neutrino simulations. The fractional calculus formulation reveals fundamental mathematical connections between neutrino flavor evolution and other systems governed by long-range temporal correlations.
I Introduction
Core-collapse supernovae (CCSNe) represent one of the most energetic and complex astrophysical phenomena in the Universe, marking the cataclysmic death of massive stars ($M\gtrsim 8M_{\odot}$). These explosions not only seed the interstellar medium with heavy elements essential for planetary formation and life, but also give birth to neutron stars and black holes [28, 12, 39]. The detection of gravitational waves from binary neutron star mergers [2, 3] and the anticipated observation of neutrinos from the next Galactic supernova by next-generation detectors such as Hyper-Kamiokande [27] and DUNE [5] underscore the importance of understanding these events through multi-messenger astronomy.
Despite decades of theoretical and computational advancements, the precise explosion mechanism of CCSNe remains an open question [29, 13, 24]. Current three-dimensional simulations suggest that the neutrino-heating mechanism, aided by hydrodynamic instabilities such as the standing accretion shock instability (SASI) and neutrino-driven convection, can successfully revive the stalled shock in many progenitor models [4, 16]. Turbulence plays a crucial role in this process by enhancing the efficiency of neutrino heating and facilitating shock revival [15, 14]. Recent high-resolution simulations reveal that turbulent energy cascades follow Kolmogorov-like scaling in the inertial range [34], but the precise characterization of turbulence in supernova environments remains challenging [40].
Neutrinos, carrying away $\sim 99\%$ of the gravitational binding energy ($\sim 3\times 10^{53}$ erg), are the primary cooling agents and potential drivers of the explosion through the neutrino-heating mechanism [28, 38]. Their flavor evolution is profoundly affected by both coherent forward scattering (the Mikheyev-Smirnov-Wolfenstein (MSW) effect) [36, 45] and collective oscillations due to neutrino-neutrino interactions [20, 19, 26]. Recent studies have identified fast flavor instabilities that can occur on centimeter scales and nanosecond timescales [8, 17, 21, 23], potentially altering the neutrino energy deposition and nucleosynthesis yields [41, 22, 44]. Moreover, muon creation in supernova matter has been shown to facilitate neutrino-driven explosions [9], adding further complexity to the flavor evolution landscape.
In addition to vacuum and matter effects, turbulence in the supernova envelope introduces stochastic density fluctuations that can significantly impact neutrino flavor evolution. The pioneering work of Loreti and Balantekin [35] established the framework for treating neutrino oscillations in noisy media, demonstrating that density fluctuations can induce decoherence and modify survival probabilities. Subsequent studies have extended this analysis to various turbulent environments [33, 43, 31, 42]. For power-law correlated turbulence, the master equation acquires a non-Markovian character with memory kernels scaling as $K(t)\propto t^{-\nu}$, where $\nu$ is the spectral index determined by the turbulence power spectrum [32, 10, 1]. Recently, Mukhopadhyay and Sen [37] investigated how turbulence signatures in supernova neutrinos could be probed in upcoming detectors, highlighting the observational relevance of this phenomenon.
The mathematical treatment of neutrino decoherence in power-law correlated turbulent media presents significant challenges due to the non-Markovian nature of the evolution equations. The master equation with memory kernel $K(t)\propto t^{-\nu}$ naturally connects to fractional calculus, where the time convolution integral represents a fractional integral operator of order $1-\nu$ [30]. For $\nu<1$, this corresponds to the Riemann-Liouville fractional integral, while for $\nu\geq 1$, regularization at small time scales becomes necessary to ensure mathematical consistency. Fractional calculus has found widespread application in describing anomalous diffusion, viscoelastic materials, and non-Markovian processes in open quantum systems [11]. In the context of neutrino physics, the emergence of Mittag-Leffler functions $E_{\alpha,\beta}(z)$ as natural solutions to fractional differential equations provides a powerful analytical framework [18].
While approximate and numerical treatments of neutrino propagation in turbulent media exist [35, 32]—ranging from white-noise models to power-law correlated fluctuations (numerical)—a complete analytical solution for the survival probability with generic power-law memory kernels has remained elusive. Previous approaches often relied on perturbative expansions or specialized limits, leaving the full mathematical structure unexplored. The power-law form of the memory kernel, $K(t)\propto t^{-\nu}$, suggests a natural connection to fractional calculus, yet this connection has not been fully exploited to obtain closed-form analytical solutions.
Several critical gaps in the current understanding warrant systematic investigation: (i) the explicit formulation of the neutrino propagation problem within a fractional calculus framework; (ii) the mathematical regularization required for spectral exponents $\nu\geq 1$ to handle ultraviolet divergences at short time scales ($t\to 0$); (iii) the derivation of exact analytical solutions for arbitrary $\nu$, expressed through special functions like Mittag-Leffler functions; (iv) the distinctive mathematical features arising for negative spectral exponents ($\nu<0$) corresponding to extremely red spectra; (v) the transition between algebraic and exponential decoherence regimes controlled by regularization; and (vi) the establishment of explicit connections between neutrino decoherence and anomalous diffusion phenomena through their shared fractional calculus structure.
This work addresses these gaps by developing a comprehensive exact theory based on fractional calculus. We establish the precise correspondence between the neutrino master equation with power-law memory and fractional differential equations. Special attention is paid to the ultraviolet divergence of the correlation function for $\nu\geq 1$, which we regularize by introducing an exponential cutoff $\exp(-\epsilon/t)$ at the dissipation time scale $\epsilon$, ensuring mathematical consistency while preserving the essential power-law behavior in the inertial range. This regularization enables the derivation of exact, closed-form solutions valid for all $\nu$.
The exact solution reveals a rich mathematical structure: for $\nu<1$, the memory integral corresponds to a Riemann-Liouville fractional integral of order $1-\nu$, while for $\nu<0$, it becomes a higher-order fractional operator. The survival probability is expressed exactly in terms of Mittag-Leffler functions, which naturally interpolate between initial algebraic decay and eventual exponential relaxation. This formulation provides not only a complete analytical description but also efficient computational methods through established properties of fractional calculus and special functions.
Our work demonstrates that the fractional calculus approach offers a unifying mathematical framework that connects neutrino decoherence in turbulent matter with broader phenomena in statistical physics, including anomalous diffusion and viscoelastic relaxation. This perspective yields new insights into the interplay between spectral index, regularization scale, and decoherence efficiency, providing both fundamental understanding and practical tools for supernova neutrino simulations.
In this work, we establish the complete mathematical formalism for the exact theory of neutrino decoherence in red turbulence. We provide a rigorous derivation of the master equation for the neutrino density matrix via an open quantum systems approach, obtaining the exact non-Markovian equation through the Nakajima–Zwanzig projection technique [11]. Furthermore, we derive the exact solution for the electron neutrino survival probability in Laplace space. For red spectral turbulence, we develop a fractional calculus framework, yielding a fractional differential equation exactly solvable in terms of Mittag-Leffler functions. This fractional formulation not only provides efficient computational methods but also uncovers profound mathematical links between neutrino flavor evolution and other physical systems characterized by power-law correlations and anomalous dynamics.
This paper is organized as follows: Section II reviews the exact Laplace-space solution with small-scale regularization. Section III develops the analytic solution via fractional calculus. Section IV summarizes our findings and outlines directions for future research. In a companion paper [7], we extend this analysis to include collective neutrino oscillations in turbulent environments.
II Exact Solution in Laplace Space
II.1 Physical Picture and Assumptions
We begin by outlining the physical scenario and the key assumptions underlying our derivation. The propagation of neutrinos through turbulent matter in core-collapse supernovae involves multiple scales:
  • Microscopic scale ($\sim 10^{-12}$ cm): Neutrino flavor oscillations governed by quantum mechanics.
  • Turbulence inertial range ($10^{2}$$10^{6}$ cm): Compressible hydrodynamic turbulence with power-law scaling [25].
  • Supernova scale ($\sim 10^{8}$ cm): Overall system size.
Our analysis bridges these scales through the following well-justified approximations:
Assumption 1: Gaussian Statistics for Density Fluctuations. We assume the turbulent density fluctuations $\delta n_{e}(\mathbf{x},t)$ form a Gaussian random field.
Assumption 2: Homogeneous, Isotropic, Stationary Turbulence. We assume statistical homogeneity (translation invariance), isotropy (rotation invariance), and stationarity (time translation invariance). These are standard assumptions in power-law turbulence theory and are reasonable for fully developed turbulence far from boundaries.
Assumption 3: Passive Scalar Approximation. We treat density fluctuations as a passive scalar advected by the velocity field but not significantly affecting it. This requires $\delta\rho/\rho\ll 1$, which holds for typical supernova conditions where density contrasts are $\sim 1\%$$5\%$.
Assumption 4: Frozen Turbulence for Relativistic Neutrinos. For ultra-relativistic neutrinos ($v_{\nu}\approx c$), we adopt Taylor’s frozen turbulence hypothesis. The justification is quantitative:
$\tau_{\text{cross}}$ $=\frac{L}{c}\sim\frac{10^{6}~\text{cm}}{3\times 10^{10}~\text{cm/s}}\sim 3\times 10^{-5}~\text{s}\,$, (1)
$\tau_{\text{eddy}}$ $=\frac{L}{v_{t}}\sim\frac{10^{6}~\text{cm}}{10^{7}~\text{cm/s}}\sim 0.1~\text{s}\,$, (2)
$\frac{\tau_{\text{cross}}}{\tau_{\text{eddy}}}$ $\sim 3\times 10^{-4}\ll 1\,.$ (3)
Thus neutrinos sample a nearly frozen turbulence pattern since the eddy time scale is much longer than the cross-time.
Assumption 5: Straight-Line Neutrino Trajectories. We assume neutrinos follow straight-line trajectories with constant velocity $\mathbf{v}_{\nu}$. This is excellent for relativistic neutrinos experiencing negligible deflection in the supernova medium.
II.2 From Turbulence Spectrum to Correlation Function
II.2.1 Obukhov-Corrsin Spectrum and Dimensionless Formulation
Let us begin by introducing dimensionless variables for a formal simplification. Let $n_{0}$ be a reference electron number density (typically the mean density $\bar{n}_{e}$), and $T_{0}$ a reference time (to be chosen as $1/\Delta_{m}$, where $\Delta_{m}$ is the matter-level splitting of neutrinos). In natural units ($\hbar=c=1$), length, time, and energy satisfy $L]=[T]=[E]^{-1}$. We define dimensionless coordinates and density fluctuations:
$\tilde{\mathbf{x}}=\frac{\mathbf{x}}{L_{0}},\quad\tilde{t}=\frac{t}{T_{0}},\quad\delta\tilde{n}_{e}(\tilde{\mathbf{x}},\tilde{t})=\frac{\delta n_{e}(\mathbf{x},t)}{n_{0}},$ (4)
where $L_{0}$ is a reference length (e.g., the integral scale $L$).
The dimensionless effective Hamiltonian for neutrino propagation is obtained by normalizing with $\Delta_{m}$:
$\tilde{H}=\frac{H}{\Delta_{m}}=\frac{1}{2}(-\cos 2\theta_{m}\sigma_{z}+\sin 2\theta_{m}\sigma_{x})+\tilde{V}(\tilde{t})\sigma_{z},$ (5)
where
[…]
where $\tilde{V}$ is the adimensional interaction potential and $G_{F}$ is the Fermi coupling constant $([G_{F}]=M^{-2})$. Here $\tilde{\mathbf{x}}_{\nu}(\tilde{t})=\mathbf{v}_{\nu}\tilde{t}$ is the dimensionless neutrino trajectory (with $v_{\nu}\approx 1$).
The correlation function of the dimensionless density fluctuations is defined as
[…]
which is itself dimensionless. For fully developed power-law turbulence, the correlation function exhibits power-law scaling in the inertial range, but requires regularization at both small and large scales. While large-scale correlations are naturally limited by the system size, the small-scale behavior requires careful treatment due to potential divergences at $\tau\to 0$.
II.2.2 Small-Scale Regularized Correlation Function
The power-law correlation $K(\tau)\propto\tau^{-\nu}$ diverges as $\tau\to 0$ for $\nu\geq 1$, representing an unphysical ultraviolet divergence. In real turbulent systems, this divergence is regularized by physical processes at the dissipation scale. We introduce a small-scale cutoff at the dissipation time scale $\tau_{\text{diss}}=l_{\text{diss}}/c$, where $l_{\text{diss}}$ is the Kolmogorov dissipation scale:
$l_{\text{diss}}\sim\left(\frac{\nu_{\text{kin}}^{3}}{\varepsilon}\right)^{1/4},$ (8)
with $\nu_{\text{kin}}$ being the kinematic viscosity and $\varepsilon$ the turbulent energy dissipation rate per unit mass.
The regularized correlation function takes the form:
[…]
For $\nu<1$, the correlation function is integrable at $\tau=0$ even without regularization, but we maintain the same form for consistency. The normalization constant $A$ is chosen such that $\int_{0}^{\infty}\tilde{\mathcal{C}}(\tilde{\tau})d\tilde{\tau}=1$ when considering the effective damping rate.
II.2.3 Exact Non-Markovian Master Equation from Nakajima–Zwanzig Projection
We now present a rigorous, non-perturbative derivation of the master equation governing neutrino decoherence in red turbulent matter using the Nakajima–Zwanzig (NZ) projection operator formalism [11]. The essential advantage of the NZ approach is that it treats the memory kernel exactly for classical Gaussian noise, automatically resumming all cumulants that contribute to the phase‑diffusion process.
Consider the total Hilbert space $\mathcal{H}_{\mathrm{total}}=\mathcal{H}_{S}\otimes\mathcal{H}_{\mathrm{env}}$, where $\mathcal{H}_{S}$ is the two‑flavor neutrino space and $\mathcal{H}_{\mathrm{env}}$ describes the turbulent matter field. The Hamiltonian is
$H=H_{0}\otimes\mathbb{I}_{\mathrm{env}}\;+\;\mathbb{I}_{S}\otimes H_{\mathrm{env}}\;+\;V\otimes\sigma_{z}\;,$ (10)
with $H_{0}=\frac{\Delta_{m}}{2}(-\cos 2\theta_{m}\sigma_{z}+\sin 2\theta_{m}\sigma_{x})$ the matter‑basis Hamiltonian and $V$ the classical stochastic potential of the density fluctuations. Because $V(t)$ is a classical (commuting) random variable, the interaction super‑operator
$\mathcal{L}_{\mathrm{int}}(t)\,\bullet\;\equiv\;V(t)\bigl[\sigma_{z},\,\bullet\,\bigr]$ (11)
satisfies $\mathcal{L}_{\mathrm{int}}(t),\mathcal{L}_{\mathrm{int}}(t^{\prime})]=0$ for all $t,t^{\prime}$. This commutativity allows the NZ memory kernel to be evaluated without any truncation of the cumulant series.
Define the projection $\mathcal{P}$ that extracts the relevant system part while averaging over the environment:
$\mathcal{P}\,\rho(t)\;\equiv\;\mathrm{Tr}_{\mathrm{env}}\bigl[\rho(t)\bigr]\otimes\rho_{\mathrm{env}}^{\mathrm{eq}}\;,$ (12)
where $\rho_{\mathrm{env}}^{\mathrm{eq}}$ is the stationary state of the turbulence; its complement is $\mathcal{Q}=\mathbb{I}-\mathcal{P}$. The exact NZ equation for the reduced density matrix $\rho_{S}(t)\equiv\mathrm{Tr}_{\mathrm{env}}\rho(t)$ reads
$\frac{d}{dt}\rho_{S}(t)\;=\;-i\mathcal{L}_{0}\,\rho_{S}(t)\;-\;\int_{0}^{t}d\tau\;\mathcal{K}(\tau)\,\rho_{S}(t-\tau)\;,$ (13)
with $\mathcal{L}_{0}\,\bullet\equiv[H_{0},\,\bullet]$ and the memory super‑kernel
$\mathcal{K}(\tau)\;=\;\mathcal{P}\,\mathcal{L}_{\mathrm{int}}\,e^{-i\mathcal{Q}\mathcal{L}\tau}\,\mathcal{Q}\,\mathcal{L}_{\mathrm{int}}\,\mathcal{P}\;.$ (14)
For a classical Gaussian field the exponential factor can be factorized exactly. Let $\Phi(\tau)=\frac{1}{2}\!\int_{0}^{\tau}\!d\tau_{1}\!\int_{0}^{\tau_{1}}\!d\tau_{2}\,\langle V(\tau_{1})V(\tau_{2})\rangle$, then
$\mathcal{K}(\tau)=\langle V(0)V(\tau)\rangle\,e^{-\Phi(\tau)[\sigma_{z},[\sigma_{z},\cdot]]}\,[\sigma_{z},[\sigma_{z},\cdot]].$ (15)
The exponential term embodies the complete resummation of the phase‑diffusion cumulants.
Inserting the regularized power‑law correlation
$\langle V(0)V(\tau)\rangle\;=\;\frac{G_{F}^{2}\,n_{0}^{2}\,A}{2\Delta_{m}^{2}}\,\tau^{-\nu}\,e^{-\epsilon/\tau}\qquad(\nu>1)\;,$ (16)
and introducing the dimensionless time $x=\Delta_{m}t$, $u=\Delta_{m}\tau$, and $\epsilon=\Delta_{m}\tau_{\mathrm{diss}}$, we obtain the exact non‑Markovian master equation
$\frac{d}{dx}\langle\rho(x)\rangle\;=\;$ $-i\bigl[\tilde{H}_{0},\,\langle\rho(x)\rangle\bigr]$
$-\;\xi\int_{0}^{x}du\;\mathcal{K}(u;\nu,\epsilon)\,\bigl[\sigma_{z},\bigl[\sigma_{z},\,\langle\rho(x-u)\rangle\bigr]\bigr]\;,$ (17)
where $\tilde{H}_{0}=H_{0}/\Delta_{m}$,
$\xi\;\equiv\;\frac{G_{F}^{2}\,n_{0}^{2}\,A}{2\Delta_{m}^{2}}\,\int_{0}^{\infty}du\;u^{-\nu}e^{-\epsilon/u}\;=\;\frac{G_{F}^{2}\,n_{0}^{2}\,A}{2\Delta_{m}^{2}}\,\epsilon^{1-\nu}\,\Gamma(\nu-1)\;,$ (18)
and the unit‑normalised kernel is
$\mathcal{K}(u;\nu,\epsilon)\;=\;\frac{u^{-\nu}e^{-\epsilon/u}}{\epsilon^{1-\nu}\,\Gamma(\nu-1)}\;.$ (19)
Equation (17) is the exact non‑Markovian master equation for neutrino propagation in red turbulence, valid non‑perturbatively for classical Gaussian noise.
The structure of the double commutator $\sigma_{z},[\sigma_{z},\cdot]]$ permits a clean separation between the diagonal (population) and off‑diagonal (coherence) components of the density matrix. To see this explicitly, write the $2\times 2$ density matrix in the high‑density matter basis as
$\rho=\begin{pmatrix}\rho_{11}&\rho_{12}\\ \rho_{21}&\rho_{22}\end{pmatrix}$, (20)
with $\rho_{21}=\rho_{12}^{*}$ for a Hermitian matrix. A direct calculation gives
$[\sigma_{z},\rho]$ $=\begin{pmatrix}0&2\rho_{12}\\ -2\rho_{21}&0\end{pmatrix}$, (21)
$[\sigma_{z},[\sigma_{z},\rho]]$ $=\begin{pmatrix}0&4\rho_{12}\\ 4\rho_{21}&0\end{pmatrix}$. (22)
Thus the double commutator acts as a projection onto the off‑diagonal entries, annihilating the diagonal ones.
In the high‑density limit $\cos 2\theta_{m}\simeq 1$, the effective Hamiltonian in the matter basis reduces to
$\tilde{H}_{0}\;=\;\frac{1}{2}\bigl(-\cos 2\theta_{m}\,\sigma_{z}+\sin 2\theta_{m}\,\sigma_{x}\bigr)\;\simeq\;-\frac{1}{2}\,\sigma_{z}.$ (23)
Its commutator with $\rho$ is simply
$\bigl[\tilde{H}_{0},\rho\bigr]\;\simeq\;-\frac{1}{2}\bigl[\sigma_{z},\rho\bigr].$ (24)
Insert these expressions into the exact NZ master equation (17). The equation for the off‑diagonal element $\rho_{12}(x)$ becomes
$\frac{d\rho_{12}}{dx}$ $=-i\bigl[\tilde{H}_{0},\rho\bigr]_{12}-\xi\!\int_{0}^{x}\!du\;\mathcal{K}(u;\nu,\epsilon)\,\bigl[\sigma_{z},[\sigma_{z},\rho(x-u)]\bigr]_{12}$
$\simeq-i\Bigl(-\frac{1}{2}\cdot 2\rho_{12}\Bigr)-\xi\!\int_{0}^{x}\!du\;\mathcal{K}(u;\nu,\epsilon)\,4\rho_{12}(x-u)$
$=i\rho_{12}(x)-4\xi\!\int_{0}^{x}\!du\;\mathcal{K}(u;\nu,\epsilon)\,\rho_{12}(x-u).$ (25)
Here we used $\tilde{H}_{0},\rho]_{12}=-\rho_{12}$ and $\sigma_{z},[\sigma_{z},\rho]]_{12}=4\rho_{12}$.
Defining $\chi_{\nu}\equiv 2\xi$ and taking the Laplace transform,
$\int_{0}^{\infty}e^{-sx}\rho_{12}(x)\,dx\equiv\widetilde{\rho}_{12}(s),$ (26)
equation (25) transforms to
$s\widetilde{\rho}_{12}(s)-\rho_{12}(0)=i\widetilde{\rho}_{12}(s)-2\chi_{\nu}\,\tilde{\mathcal{K}}(s;\nu,\epsilon)\,\widetilde{\rho}_{12}(s),$ (27)
where $\tilde{\mathcal{K}}(s;\nu,\epsilon)$ is the Laplace transform of the normalized kernel (19). Rearranging,
$\bigl[s-i+2\chi_{\nu}\tilde{\mathcal{K}}(s;\nu,\epsilon)\bigr]\,\widetilde{\rho}_{12}(s)\;=\;\rho_{12}(0).$ (28)
For notational clarity, we define the effective damping coefficient
$\kappa_{\nu}\equiv 4\xi,$ (29)
which absorbs the numerical factors arising from the double commutator structure. The off‑diagonal equation then takes the compact form
$\boxed{\;\bigl[s+\kappa_{\nu}\tilde{\mathcal{K}}(s;\nu,\epsilon)-i\bigr]\,\widetilde{\rho}_{12}(s)\;=\;\rho_{12}(0)\;.\;}$ (30)
Here $\kappa_{\nu}=4\xi$ represents the effective turbulence strength for decoherence, with $\xi$ given by Eq. (18).
The survival probability for an electron neutrino created at $x=0$ is
$P_{ee}(x)=\frac{1}{2}+\mathrm{Re}\!\bigl[\rho_{12}(x)\,e^{-ix}\bigr],$ (31)
where the phase $e^{-ix}$ originates from the matter‑basis oscillation factor $\exp(-i\Delta_{m}t)=\exp(-ix)$ in the high‑density limit. Taking the Laplace transform and using (30),
$\tilde{P}_{ee}(s)$ $=\frac{1}{2}\int_{0}^{\infty}e^{-sx}\,dx$
$+\frac{1}{2}\int_{0}^{\infty}e^{-sx}\,\bigl[\rho_{12}(x)e^{-ix}+\rho_{12}^{*}(x)e^{+ix}\bigr]dx$
$=\frac{1}{2}\cdot\frac{1}{s}+\frac{1}{2}\Bigl[\frac{\rho_{12}(0)}{s+\kappa_{\nu}\tilde{\mathcal{K}}(s)-i}+\frac{\rho_{12}^{*}(0)}{s+\kappa_{\nu}\tilde{\mathcal{K}}(s)+i}\Bigr].$ (32)
For an initial electron neutrino, $\rho_{12}(0)=\frac{1}{2}\sin 2\theta_{m}\simeq 0$ in the high‑density limit, but retaining the formal expression and setting $\rho_{12}(0)=\rho_{12}^{*}(0)=1/2$ (which amounts to an overall normalization) gives the compact result
$\tilde{P}_{ee}(s)\;=\;\frac{1}{2}\Bigl[\frac{1}{s}+\frac{1}{s+\kappa_{\nu}\tilde{\mathcal{K}}(s;\nu,\epsilon)-i}+\frac{1}{s+\kappa_{\nu}\tilde{\mathcal{K}}(s;\nu,\epsilon)+i}\Bigr]\;.\;$} (33)
Combining the three terms yields the all‑orders Laplace‑space expression quoted in equation (34).
Combining the three terms yields the compact, all‑orders expression
$\boxed{\;\tilde{P}_{ee}(s)\;=\;\frac{s\bigl[s+\kappa_{\nu}\tilde{\mathcal{K}}(s;\nu,\epsilon)\bigr]+1}{s\bigl\{\bigl[s+\kappa_{\nu}\tilde{\mathcal{K}}(s;\nu,\epsilon)\bigr]^{2}+1\bigr\}}\;.\;}$ (34)
This result, which emerges naturally from the non‑perturbative NZ treatment, encapsulates the complete interplay between coherent oscillations (poles at $s+\kappa_{\nu}\tilde{\mathcal{K}}(s)=\pm i$) and the algebraic decoherence induced by the long‑range correlations of red turbulence.
II.3 Relation to Fractional Calculus
The memory integral with power-law kernel exhibits a natural connection to fractional calculus for the physically most relevant case of red noise ($\nu<1$).
Fractional integral representation for $\nu<1$.
For the Riemann-Liouville fractional integral of order $\alpha>0$ [6],
$I^{\alpha}[f](x)=\frac{1}{\Gamma(\alpha)}\int_{0}^{x}(x-u)^{\alpha-1}f(u)\,du.$ (35)
Changing integration variable w=xuw=x-u in our memory term gives
$\int_{0}^{x}u^{-\nu}f(x-u)\,du=\int_{0}^{x}(x-w)^{-\nu}f(w)\,dw.$ (36)
For $\nu<1$, this is proportional to a fractional integral:
$\int_{0}^{x}u^{-\nu}f(x-u)\,du=\Gamma(1-\nu)\,I^{1-\nu}[f](x),$ (37)
where the convergence of the integral at $u=0$ is guaranteed by $\nu<1$.
Thus, for red noise spectra ($\nu<1$), the non-Markovian master equation incorporates a Riemann-Liouville fractional integral operator of order $1-\nu$. This fractional structure explains the emergence of Mittag-Leffler functions in the solution, which interpolate between initial power-law decay and eventual exponential relaxation.
Regularization for the general case.
For completeness, we note that the small-scale regularization $\exp(-\epsilon/u)$ ensures mathematical consistency for any $\nu$. However, in realistic turbulent environments, spectral indices are typically in the range $0<\nu\leq 1$, corresponding to red noise spectra where low-frequency fluctuations dominate. The case $\nu>1$ (blue noise) is physically less relevant for turbulence in supernovae, as it would imply energy concentrating at high frequencies, contrary to the observed inertial-range scaling of turbulent cascades.
The fractional calculus connection provides not only mathematical elegance but also practical computational tools, as fractional differential equations with Mittag-Leffler function solutions are well-studied in the theory of anomalous diffusion and non-Markovian processes.
III Analytic Solution via Fractional Calculus
For $\nu<0$, the regularized correlation function exhibits a distinctive mathematical structure:
$K(t)=\xi\,t^{|\nu|}\exp\left(-\frac{\epsilon}{t}\right),\quad|\nu|=-\nu>0,$ (38)
where we have used the same regularization parameter $\epsilon$ as in Eq. (9) for consistency.
The Laplace transform yields:
$\tilde{K}_{\nu}(z)=\xi\,\Gamma(1-\nu)\,(z+\gamma)^{\nu-1},\quad\gamma=\frac{1}{\tau_{c}}.$ (39)
The key insight: The memory integral in the master equation becomes a fractional integral operator of order $|\nu|+1$.
From the definition of Riemann-Liouville fractional integral in Eq.35, our Master equation can be elegantly written as [6]:
$\frac{d}{dt}\rho(t)=-i[H_{0},\rho(t)]-\xi\Gamma(1-\nu)\,I^{|\nu|+1}[\sigma_{z},[\sigma_{z},\rho]].$ (40)
For $\nu=-1$ ($|\nu|=1$), this is a regular double integral. For $\nu=-2$ ($|\nu|=2$), it’s a triple integral.
III.1 Exact Solution via Operator Method
Consider the projection onto flavor coherence. Define $Q(t)=\langle\nu_{e}|\rho(t)|\nu_{x}\rangle$ in the matter basis. The equation simplifies to:
$\frac{d}{dt}Q(t)$ $=i\Delta_{m}Q(t)-4\xi\Gamma(1-\nu)I^{|\nu|+1}[Q](t),$ (41)
$Q(0)$ $=\frac{1}{2}\sin 2\theta_{m}\,.$ (42)
Applying the Laplace transform, we obtain
$(s-i\Delta_{m})\tilde{Q}(s)-Q(0)=-4\xi\Gamma(1-\nu)\frac{\tilde{Q}(s)}{s^{|\nu|+1}}\,.$ (43)
Thus:
$\tilde{Q}(s)=\frac{Q(0)}{s-i\Delta_{m}+4\xi\Gamma(1-\nu)s^{-(|\nu|+1)}}.$ (44)
III.2 Analytic Structure for Integer $|\nu|$
For $|\nu|=m\in\mathbb{Z}^{+}$, multiply both sides by $s^{m+1}$:
$\big[s^{m+2}-i\Delta_{m}s^{m+1}+4\xi\,m!\big]\tilde{Q}(s)\;=\;s^{m+1}Q(0).$ (45)
Inverse transforming yields an $(m+2)$-th order differential equation:
$Q^{(m+2)}(t)-i\Delta_{m}Q^{(m+1)}(t)-4\xi\Gamma(1+m)Q(t)=0.$ (46)
This is the key simplification: The nonlocal memory integral transforms into a local differential operator!
III.3 Mittag-Leffler Function Representation
The exact solution of the fractional master equation (40) can be expressed in terms of multivariate Mittag-Leffler functions. For the off-diagonal element $Q(t)$ in the high-density limit, one finds
$Q(t)=Q(0)\,E_{(|\nu|+2,\;|\nu|+1)}\Bigl(i\Delta_{m}t^{|\nu|+2},\;-4\xi\,\Gamma(1+|\nu|)\,t^{|\nu|+2}\Bigr),$ (47)
where the $(p,q)$-variate Mittag-Leffler function is defined by the double series
$E_{(p,q)}(a,b)\;=\;\sum_{n=0}^{\infty}\frac{a^{n}t^{pn}}{(pn)!}\;{}_{1}F_{1}\!\left(\begin{array}[]{c}1\\ qn+1\end{array};bt^{q}\right),$ (48)
with $_{1}F_{1}$ being the confluent hypergeometric function. This representation reveals the interplay between oscillatory dynamics (governed by the parameter $a=i\Delta_{m}$) and algebraic damping (governed by $b=-4\xi\,\Gamma(1+|\nu|)$).
III.4 Physical Interpretation: Accelerated Decoherence
The fractional integral operator $I^{|\nu|+1}$ represents a smearing of quantum phase information over an algebraically weighted past. This leads to:
  1. 1.
    Enhanced memory effects: The kernel $t^{|\nu|}$ gives greater weight to recent history compared to white noise.
  2. 2.
    Algebraic-to-exponential crossover: At $t\sim\tau_{c}$, the exponential cutoff $e^{-t/\tau_{c}}$ dominates, recovering Markovian behavior.
  3. 3.
    Spectral reddening: The Fourier transform $\tilde{K}_{\nu}(\omega)\sim|\omega|^{-(|\nu|+1)}$ shows enhanced low-frequency power.
IV Conclusion
In this work, we have established a comprehensive exact mathematical framework for analyzing neutrino decoherence in power-law correlated turbulent matter, as encountered in core-collapse supernovae. Our approach bridges the theoretical foundations of fractional calculus with practical applications in neutrino astrophysics, providing rigorous analytical solutions that capture the full non-Markovian dynamics of flavor evolution.
The central achievement of this work is the development of a physically motivated regularization scheme for the ultraviolet divergence that arises in power-law correlation functions when the spectral index $\nu\geq 1$. By introducing an exponential cutoff $\exp(-\epsilon/t)$ at the dissipation time scale $\epsilon$, we ensure mathematical consistency of all integrals and operators while preserving the essential power-law behavior in the inertial range. This regularization enables a well-defined fractional calculus representation of the exact non-Markovian master equation, obtained through the Nakajima-Zwanzig projection technique for classical Gaussian noise.
We derive the exact Laplace-space solution for the neutrino survival probability, expressed in the compact form, and obtain time-domain solutions analytically expressed through Mittag-Leffler functions $E_{\alpha,\beta}(z)$. These special functions naturally emerge as the fundamental building blocks of the solution, revealing the intrinsic mathematical structure of the decoherence process.
Our analysis elucidates the profound connection between neutrino flavor evolution and fractional calculus. For red turbulence spectra ($\nu<1$), the memory kernel corresponds to a Riemann-Liouville fractional integral of order $1-\nu$, while for extremely red spectra ($\nu<0$) it becomes a higher-order fractional operator. This fractional structure explains the characteristic non-exponential decoherence, with an initial power-law decay $P_{ee}(x)\sim 1-C_{\nu}x^{2-\nu}$ transitioning to exponential damping at late times.
We provided the first detailed study of neutrino decoherence for negative spectral exponents ($\nu<0$), corresponding to extremely red turbulence spectra. In these cases, the memory integral translates into higher-order fractional operators, leading to distinct dynamical scaling and mathematical features. The exact solutions for these regimes demonstrate how spectral reddening affects decoherence efficiency and memory effects.
The physical implications for supernova neutrinos are significant. Our exact analytical results serve as a crucial benchmark for numerical simulations and provide clear guidelines: (i) the high-density matter basis should be used when $\Delta_{m}\gg\Delta_{v}$ to properly account for matter effects; (ii) the small-scale cutoff must be included for spectral indices $\nu\geq 1$ to ensure mathematical rigor; (iii) the fractional calculus formulation offers efficient computational methods through established properties of Mittag-Leffler functions; and (iv) the interplay between spectral index, regularization scale, and decoherence efficiency can now be systematically studied within our exact framework.
The theoretical connections established in this work are noteworthy. Neutrino propagation in turbulent matter provides a physically rich, exactly solvable example of non-Markovian dynamics with calculable memory effects. The parallels with anomalous diffusion processes—both described by fractional differential equations and Mittag-Leffler functions—suggest potential cross-fertilization between neutrino physics and other fields studying systems with long-range temporal correlations. Our fractional calculus approach offers a unifying mathematical framework that connects neutrino decoherence with broader phenomena in statistical physics, including viscoelastic relaxation and subdiffusive transport.
Several promising directions for future research emerge from this work. These include: (i) extending the analysis to multi-dimensional anisotropic turbulence with realistic supernova profiles; (ii) incorporating turbulence effects into the nonlinear evolution equations for collective neutrino oscillations; (iii) developing efficient numerical algorithms based on our exact fractional calculus framework for implementation in multi-dimensional supernova simulations; (iv) establishing clearer connections between the spectral index $\nu$ and fundamental parameters of hydrodynamic turbulence through direct numerical simulations; (v) exploring observational consequences for neutrino signals from the next Galactic supernova, potentially using turbulence signatures as probes of supernova dynamics; and (vi) applying the fractional calculus approach to other astrophysical systems where stochastic fluctuations affect quantum coherence.
In summary, we have presented a complete exact mathematical framework for analyzing neutrino decoherence in power-law correlated turbulent environments. By combining rigorous regularization techniques with fractional calculus methods, we have developed closed-form analytical solutions that capture the full interplay between coherent oscillations and algebraic decoherence induced by long-range temporal correlations. Our work establishes firm mathematical foundations for future studies of neutrino-turbulence interactions in astrophysical environments, particularly in the complex conditions of core-collapse supernovae. The fractional calculus perspective not only provides practical computational tools but also reveals deep connections between neutrino flavor evolution and other physical systems exhibiting anomalous dynamics. As supernova simulations continue to increase in sophistication and next-generation neutrino detectors come online, the exact analytical treatment of turbulence effects developed here will prove essential for accurate interpretation of neutrino signals as probes of extreme astrophysical phenomena.
Acknowledgements.
We thank Thierry Foglizzo for insightful discussions and the developers of the mpmath Python library for high-precision numerical evaluation of special functions. During the preparation of this work, the authors used DeepSeekV3 to improve readability and language. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication. This work is supported by the Fundamental Research Funds for the Central Universities under No. 020114380057, and by K. C. Wong Educational Foundation. A.A. is supported by the National Science Foundation of China (NSFC) through the grant No. 12350410358; the Talent Scientific Research Program of College of Physics, Sichuan University, Grant No. 1082204112427; the Fostering Program in Disciplines Possessing Novel Features for Natural Science of Sichuan University, Grant No.2020SCUNL209 and 1000 Talent program of Sichuan province 2021.